74 degrees is the same as ___ radians. Rounded to the nearest hundredth of a radian?
1
Expert's answer
2019-05-20T02:58:40-0400
In a circle, 2π radians correspond to 360 degrees. Then we write the following equation

"{x \over \pi} ={ 74^0 \over 180^0}; \rArr
{x=\pi \cdot \ {74^0 \over 180^0}}; \rArr x \approx 1.29 rad"

74 degrees is the same as _1.29__ radians.